# From Isolated Conversations to Hierarchical Schemas: Dynamic Tree Memory Representation for LLMs

### v2

@@ C.2 Failure Examples
| 1 | 0.529 | 7 | Level 1: 462 Level 2: 464 Level 3: 473 | The recent court testimonies have shed light on the intricate and problematic relationship between FTX and Alameda Research …, exacerbated by the entangled operations with Alameda Research and the misleading assurances given to investors and employees. |
| 2 | 0.515 | 7 | Level 1: 462 Level 2: 464 Level 3: 473 | The recent court proceedings have highlighted significant misconduct and fraudulent practices within FTX and Alameda Research … preferential treatment, which ultimately led to the collapse of FTX. |
| 3 | 0.510 | 6 | Level 1: 462 Level 2: 464 Level 3: 473 | The recent court proceedings have shed light on the intricate and fraudulent operations at FTX and Alameda Research. Gary Wang, co-founder and CTO, revealed that FTX misrepresented the amount in … Bankman-Fried at the center of the fraudulent activities. |
| 4 | 0.503 | 4 | Level 1: 462 Level 2: 463 Level 3: 475 | Sam Bankman-Fried was a terrible boyfriend Source: The Verge, Date: 2023-10-10 Bankman-Fried , … in bubbles along the main text Ellison wrote she was worried about “both actual leverage and presenting on our balance sheet.” Bankman-Fried responded with a note: “Yup, and could also get worse." |
| 5 | 0.502 | 5 | Level 1: 40 Level 2: 42 Level 3: 43 | Sam Bankman-Fried, former CEO of FTX, testified about his limited knowledge and involvement in the financial operations … at the financial mismanagement and liabilities, highlighting his insufficient oversight that led to the companies’ problems. |
| 6 | 0.499 | 6 | Level 1: 40 Level 2: 42 Level 3: 43 | The jury finally hears from Sam Bankman-Fried Source: The Verge, Date: 2023-10-28 Risk is an inherent part of a futures exchange… Alameda Research’s liabilities had gotten too high, and FTX was spending too much money on marketing. |
| 7 | 0.496 | 8 | Level 1: 462 Level 2: 464 Level 3: 473 | Today the FTX jury suffered through a code review Source: The Verge, Date: 2023-10-06 In the process of liquidating, FTX … billion in customer money was gone, but on November 7th, Bankman-Fried tweeted “FTX is fine. Assets are fine." |
| 8 | 0.496 | 5 | Level 1: 462 Level 2: 464 Level 3: 474 | In the end, the FTX trial was about the friends screwed along the way Source: The Verge, Date: 2023-10-26 The defense … liability that Ellison, Singh, Wang, and Bankman-Fried had known about in June 2022, we saw that $1.2 billion was a loan repayment to crypto lender Genesis. |
| 9 | 0.495 | 3 | Level 1: 462 Level 2: 463 Level 3: 475 | Sam Bankman-Fried and Caroline Ellison’s intertwined personal and professional relationships significantly influenced the operations of FTX and Alameda Research … eventual scrutiny of both entities. |
| 10 | 0.493 | 5 | Level 1: 462 Level 2: 464 Level 3: 939 | The trial of Sam Bankman-Fried, founder of the collapsed cryptocurrency exchange FTX, has exposed significant … of the entanglements between FTX and Alameda Research, emphasizing the fraudulent nature of their operations. |

### v3

@@ C.2 Failure Examples
1 | 0.529 | 7 | Level 1: 462 Level 2: 464 Level 3: 473 | The recent court testimonies have shed light on the intricate and problematic relationship between FTX and Alameda Research …, exacerbated by the entangled operations with Alameda Research and the misleading assurances given to investors and employees. |
2 | 0.515 | 7 | Level 1: 462 Level 2: 464 Level 3: 473 | The recent court proceedings have highlighted significant misconduct and fraudulent practices within FTX and Alameda Research … preferential treatment, which ultimately led to the collapse of FTX. |
3 | 0.510 | 6 | Level 1: 462 Level 2: 464 Level 3: 473 | The recent court proceedings have shed light on the intricate and fraudulent operations at FTX and Alameda Research. Gary Wang, co-founder and CTO, revealed that FTX misrepresented the amount in … Bankman-Fried at the center of the fraudulent activities. |
4 | 0.503 | 4 | Level 1: 462 Level 2: 463 Level 3: 475 | Sam Bankman-Fried was a terrible boyfriend Source: The Verge, Date: 2023-10-10 Bankman-Fried , … in bubbles along the main text Ellison wrote she was worried about “both actual leverage and presenting on our balance sheet.” Bankman-Fried responded with a note: “Yup, and could also get worse." |
5 | 0.502 | 5 | Level 1: 40 Level 2: 42 Level 3: 43 | Sam Bankman-Fried, former CEO of FTX, testified about his limited knowledge and involvement in the financial operations … at the financial mismanagement and liabilities, highlighting his insufficient oversight that led to the companies’ problems. |
6 | 0.499 | 6 | Level 1: 40 Level 2: 42 Level 3: 43 | The jury finally hears from Sam Bankman-Fried Source: The Verge, Date: 2023-10-28 Risk is an inherent part of a futures exchange… Alameda Research’s liabilities had gotten too high, and FTX was spending too much money on marketing. |
7 | 0.496 | 8 | Level 1: 462 Level 2: 464 Level 3: 473 | Today the FTX jury suffered through a code review Source: The Verge, Date: 2023-10-06 In the process of liquidating, FTX … billion in customer money was gone, but on November 7th, Bankman-Fried tweeted “FTX is fine. Assets are fine." |
8 | 0.496 | 5 | Level 1: 462 Level 2: 464 Level 3: 474 | In the end, the FTX trial was about the friends screwed along the way Source: The Verge, Date: 2023-10-26 The defense … liability that Ellison, Singh, Wang, and Bankman-Fried had known about in June 2022, we saw that $1.2 billion was a loan repayment to crypto lender Genesis. |
9 | 0.495 | 3 | Level 1: 462 Level 2: 463 Level 3: 475 | Sam Bankman-Fried and Caroline Ellison’s intertwined personal and professional relationships significantly influenced the operations of FTX and Alameda Research … eventual scrutiny of both entities. |
10 | 0.493 | 5 | Level 1: 462 Level 2: 464 Level 3: 939 | The trial of Sam Bankman-Fried, founder of the collapsed cryptocurrency exchange FTX, has exposed significant … of the entanglements between FTX and Alameda Research, emphasizing the fraudulent nature of their operations. |